A parallel plate capacitor is made of two rectangular plates of area 1.0 m^2 and the
separation between plates is 2.0 mm. A voltage of 100 V is applied across the
plates. Calculate the capacitance of the capacitor if a dielectric material of dielectric
constant 3.0 is introduced between the plates. What is the charge stored on each
plate of the capacitor?
1
Expert's answer
2019-03-22T10:37:01-0400

The expression for the capacitance of a plane capator can be written as:


"C = \frac{\varepsilon \varepsilon_0 S}{d},"

where S is the area of the plates, d is the separation between the plates, epsilon is the dielectric material constant, epsilon0 is the electric constant.


Substituting the numerical values, we obtain:


"C=\frac{8.85 \cdot 10^{-12} \cdot 3 \cdot 1}{2 \cdot 10^{-3}} \approx 13.3 \cdot 10^{-9} \, F = 13.3 \, nF"

The charge stored on each plate can be calculated as:


"q = CV"

Substituting the numerical values, we obtain:


"q = 13.3 \cdot 10^{-9} \cdot 10^{2} = 13.3 \cdot 10^{-7} \, C \approx 1.3 \, \mu C"

Answer: 13.3 nF, 1.3 muC.